OpenStudy (anonymous):

If an object is dropped from a height of 200 feet, the function h(t)=-16t^2+200 gives the height of the object after t seconds. Approximately, when will the object hit the ground? A.200.00 Seconds B.184.00 Seconds C.3.54 Seconds D. 0.78 seconds

OpenStudy (anonymous):

Really need help!

OpenStudy (anonymous):

The object hits the ground at time \(t\) that makes the height zero. Solve for \(t\): \[-16t^2+200=0\]

OpenStudy (anonymous):

No clue man.

OpenStudy (anonymous):

\[16t^2=200~~\iff~~t^2=\frac{200}{16}\]

OpenStudy (anonymous):

Problem Is in not very smart.

OpenStudy (anonymous):

And I'm not a particularly patient person.

OpenStudy (anonymous):

You should review your algebra.